OpenStudy (anonymous):

1+ cos^2x - (1-cos^2x) becomes 2 cos^2x. I thought it would equal 0. Is it because cos is an even function ???

OpenStudy (ajprincess):

1+cos^2x-(1-cos^2x)=1+cos^2x-1+cos^2x=2cos^2x (-)*(-)=(+)

OpenStudy (ajprincess):

Does that help? @pitanje

OpenStudy (anonymous):

so the second minus becomes positive and has nothing to do with being an odd function?

OpenStudy (ajprincess):

ya

OpenStudy (ajprincess):

when u multiply a minus with a minus u get positive

OpenStudy (anonymous):

ok thanks

OpenStudy (ajprincess):

welcome:)